OpenStudy (anonymous):

A Medium Apple Has 72 Calories . This Is 3/4 Of The Calories In A Banana . How Many Calories Does A Medium Banana Have ?

OpenStudy (anonymous):

72 * 4/3 = 96 calories

OpenStudy (anonymous):

ThankYouuu '

OpenStudy (anonymous):

MedApple = 3/4 MedBanana 72 = 3/4 MedBanana 72 * 4/3 = Med Banana = 96 just a tiny bit of work, just in case you wanted or needed it.